Mona's father is thrice as old as mona after12 years she will be twice the age ofbhis daughter . Find the present
Answer by ankor@dixie-net.com(22740)

let m = Mona's age
let f = father's age
:
Write an equation for each statement, simplify as much as possible:
:
Mona's father is thrice as old as Mona
f = 3m
after 12 years she will be twice the age of his daughter.
f + 12 = 2(m + 12)
f + 12 = 2m + 24
f = 2m + 24 - 12
f = 2m + 12
Replace f with 3m, from the 1st statement
3m = 2m + 12
3m - 2m = 12
m = 12 yrs is Mona's age
I'll let you find father's age from this, check your solutions in the 2nd statement.
Find their present ages